Tommy eats only green eggs and ham. He has an income of $36. Green eggs have a price of $2 and ham has a price of $6.
a. On the graph below, draw in Tommy’s budget constraint. Put ham on the horizontal axis. Suppose Tommy’s income increases to $54. Also suppose that the price of ham and egg each rise by $10. Construct Tommy’s new budget constraint. What is the difference between the new and old budget constraints?
1
Expert's answer
2019-04-15T10:32:12-0400

Tommy’s budget constraint is a line, which begins with 36/2 = 18 units of green eggs on y-axis and 36/6 = 6 units of ham on x-axis.

If Tommy’s income increases to $54 and the price of ham and egg each rise by $10, then Tommy’s budget constraint will shift and will begin with 54/(2 + 10) = 4.5 units of green eggs on y-axis and 54/(6 + 10) = 3.375 units of ham on x-axis. So, the new budget constraint will shift to the left comparing to the old one.